For the Ring, R= Z2[x]/<x^8-1>. Find zero divisors and nilpotent elements if any.
1
Expert's answer
2020-09-28T13:29:37-0400

We know that , If "\mathbb{F}" is a finite field , then

"\frac{ \mathbb{F}[x]}{<f(x)>}" would be a finite dimensional vector space over a finite field and thus would be a finite set .

Now "\frac{\mathbb{F}[x]}{<f(x)>}" is a domain (has no zero divisiors)

"\iff <f(x)>" is a prime ideal

"\iff f(x)=0 \ or \ f(x)" is a prime elements of the ring (Because we are in a commutative PID)

"\iff f(x)" is irreducible

(Because in a PID prime elements are same as irreducible)

Now in your case "f(x)=x^8-1" and "\mathbb{F}=\Z_2"

Therefore "f(x)" is not irreducible because 1 is a root of "f(x)" .

Now under mod 2 , "x^8-1=x^8+1"

Now "x^8+1=(x^4)^2+2x^4+1^2"

"=(x^4+1)^2"

Again, "x^4+1=(x^2)^2+2x^2+1^2"

"=(x^2+1)^2"

But ,"x^2+1=x^2+2x+1^2"

"=(x+1)^2"

"\therefore x^8+1=(x+1)^8=(x-1)^8"

........................(1)

A nonzero element "a" of a ring "R" is said to be zero divisors if there exist a nonzero element "b\in R" such that

either "ab=0 \ or \ ba=0" .

In our case , a zero divisor should be divisior of "x^8-1" and not equal to "x^8-1" .

Therefore the zero divisor of

"\frac{ \Z_2[x]}{<x^8-1>}" are "(x-1)+<x^8-1>,(x-1)^2+<x^8-1>"

,"(x-1)^3+<x^8-1>,(x-1)^4+<x^8-1>"

,

"(x-1)^5+<x^8-1>,(x-1)^6+<x^8-1>"

,"(x-1)^7+<x^8-1>"

An element "r(x)+<x^8-1>" is nilpotent if and only if there exist a "n\in \Z_{>0}" such that "(x^8-1)| r^n(x)" but "x^8-1" does not divide "r(x)."

Therefore the nilpotent element are all those element of "\frac{\Z_2[x]}{<x^8-1>}" which have zero divisor as a factor.